The equation of exchange is

an abstraction
a theory.
a, b, and c
a hypothesis.
an identity.
Answer #1

option b. i.e. a theory

The equation of exchange shows the relationship between money and price level and between money and nominal GDP which says that M * V = P * Y and it says that total sales = total spending and is a theory of macroeconomy